OpenStudy (anonymous):

Consider a small car of mass 1296 kg and a large sport utility vehicle (SUV) of mass 3640 kg. The SUV is traveling at the speed limit (v = 34 m/s). The driver of the small car travels so as to have the same kinetic energy as the SUV. Find the speed of the small car

OpenStudy (roadjester):

\(K=\dfrac 1 2 mv^2\) So let's use subscript "c" for car and "s" for SUV \(K_c=\dfrac 1 2 m_cv_c^2\) \(K_s=\dfrac 1 2 m_sv_s^2\) You want them to have the same kinetic energy so just substitute and do the math. =)

OpenStudy (anonymous):

Thank you the answer is 57 m/s

OpenStudy (roadjester):

cool